The temperature coefficient of resistance for carbon at 0oC is -0.0048/oC. What is the significance of the minus sign?
A carbon resistor has a resistance of 500Ω.Determine its resistance at 50oC
1
Expert's answer
2020-05-04T13:12:56-0400

The minus sign means that as the temperature increases, the resistance decreases.

"TCR=\frac{R_2-R_1}{R_1\times(T_2-T_1)};\\R_2-R_1=TRC\times R_1\times(T_2-T_1);\\R_2=R_1\times(TRC\times(T_2-T_1)+1);\\R_2=500\times(-0.0048\times(50-25)+1)=440\Omega;"

Answer: 1.The minus sign means that as the temperature increases, the resistance decreases.

"R_2=440\Omega"